OpenStudy (anonymous):

what does it mean when "It says Multiply. Write your answer in lowest terms and as whole or mixed numbers

OpenStudy (anonymous):

it means to multiply and then write the answer in the lowest temrs as whole or mixed numbers

OpenStudy (lilai3):

it means that ur suppose to write the fraction in simplst form..i believe

OpenStudy (anonymous):

16/15x10/48 how do i start?

OpenStudy (faman39):

cross multiply

OpenStudy (faman39):

i guess

OpenStudy (anonymous):

no its fractions

OpenStudy (faman39):

ya i kno, it kinda multiplication fraction

OpenStudy (lilai3):

u cross multiply, but first, find a common # dat can b divided by each other (they haf to be across from each other)

OpenStudy (anonymous):

can u draw out the example

OpenStudy (anonymous):

thats exactly what i said

OpenStudy (faman39):

lilai is rite

OpenStudy (anonymous):

please i need to see an example, dow it out please

OpenStudy (faman39):

kk too

OpenStudy (anonymous):

16/15 x 10/48

OpenStudy (anonymous):

?

OpenStudy (katrinakaif):

You multiply straight across Vils

OpenStudy (mani_jha):

 16*3=48 So, 16/48=1/3 So, instead of 16/48, write 1/3. Then, 10=5*2 15=5*3 So, 10/15=2/3. So just put 2/3 instead of 10/15, Multiply straight whatever is left. So, (1/3)*(2/3)=2/9 Got it?